OpenStudy (anonymous):

What is the area of a square with a side of 2.5?

OpenStudy (anonymous):

use the formula A=s^2

OpenStudy (anonymous):

use that ^

OpenStudy (anonymous):

A=(2.5)^2 A=6.25

OpenStudy (anonymous):

a = s^2 a = 2.5*2.5 a = 6.25